# Multi-LoRA Composition for Image Generation

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) is extensively utilized in text-to-image models for the accurate rendition of specific elements like distinct characters or unique styles in generated images. Nonetheless, existing methods face challenges in effectively composing multiple LoRAs, especially as the number of LoRAs to be integrated grows, thus hindering the creation of complex imagery. In this paper, we study multi-LoRA composition through a decoding-centric perspective. We present two training-free methods: LoRA Switch, which alternates between different LoRAs at each denoising step, and LoRA Composite, which simultaneously incorporates all LoRAs to guide more cohesive image synthesis. To evaluate the proposed approaches, we establish ComposLoRA, a new comprehensive testbed as part of this research. It features a diverse range of LoRA categories with 480 composition sets. Utilizing an evaluation framework based on GPT-4V, our findings demonstrate a clear improvement in performance with our methods over the prevalent baseline, particularly evident when increasing the number of LoRAs in a composition.
Machine Learning, ICML
@@ 1 Introduction
In the dynamic realm of generative text-to-image models (Ho et al., 2020; Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; Ruiz et al., 2023; Sohn et al., 2023), the integration of Low-Rank Adaptation (LoRA) (Hu et al., 2022) stands out for its ability to fine-tune image synthesis with remarkable precision and minimal computational load. LoRA excels by specializing in one element — such as a specific character, a particular clothing, a unique style, or other distinct visual aspects — and being trained to produce diverse and accurate renditions of this element in generated images. For instance, users could customize their LoRA model to generate various images of themselves, achieving an array of personalized and realistic representations. The application of LoRA not only showcases its adaptability and precision in image generation but also opens new avenues in customized digital content creation, revolutionizing how users interact with and utilize generative text-to-image models for creating tailored visual content.
@@
However, an image typically embodies a mosaic of various elements, making compositionality key to controllable image generation (Tenenbaum, 2018; Huang et al., 2023b). In pursuit of this, the strategy of composing multiple LoRAs, each focused on a distinct element, emerges as a feasible approach for advanced customization. This technique enables the digitization of complex scenes, such as virtual try-ons, merging users with clothing in a realistic fashion, or urban landscapes where users interact with meticulously designed city elements. Prior investigations into multi-LoRA compositions have explored the context of pre-trained language models (Zhang et al., 2023a; Huang et al., 2023a) or stable diffusion models (Ryu, 2023; Shah et al., 2023). These studies aim to merge multiple LoRA models to synthesize a new LoRA model by training coefficient matrices (Huang et al., 2023a; Shah et al., 2023) or through the direct addition or subtraction of LoRA weights (Ryu, 2023; Zhang et al., 2023a). Nevertheless, these approaches centered on weight manipulation could destabilize the merging process as the number of LoRAs grows (Huang et al., 2023a) and also overlook the interaction between LoRA models and base models. This oversight becomes particularly critical in diffusion models, which depend on sequential denoising steps for image generation. Ignoring the interplay between LoRAs and these steps can result in misalignments in the generative process, as shown in Figure 1, where a merged LoRA model fails to preserve the full complexity of all desired elements, leading to distorted or unrealistic images.
@@
Figure 1: Multi-LoRA composition techniques effectively blend different elements such as characters, clothing, and objects into a cohesive image. Unlike the conventional LoRA Merge approach, which can lead to detail loss and image distortion as more LoRAs are added, our methods retain the accuracy of each element and the overall image quality.
@@
In this paper, we delve into multi-LoRA composition from a decoding-centric perspective, keeping all LoRA weights intact. We present two learning-free approaches that utilize either one or all LoRAs at each decoding step to facilitate compositional image synthesis. Our first approach, LoRA Switch, operates by selectively activating a single LoRA during each denoising step, with a rotation among multiple LoRAs throughout the generation process. For instance, in a virtual try-on scenario, LoRA Switch alternates between a character LoRA and a clothing LoRA at successive denoising steps, thereby ensuring that each element is rendered with precision and clarity. In parallel, we propose LoRA Composite, a technique that draws inspiration from classifier-free guidance (Ho & Salimans, 2022). It involves calculating unconditional and conditional score estimates derived from each respective LoRA at every denoising step. These scores are then averaged to provide balanced guidance for image generation, ensuring a comprehensive incorporation of all elements. Furthermore, by bypassing the manipulation on the weight matrix but directly influencing the diffusion process, both methods allow for the integration of any number of LoRAs and overcome the limitations of recent studies that typically merge only two LoRAs (Shah et al., 2023).
@@
Experimentally, we introduce ComposLoRA, the first testbed specifically designed for LoRA-based composable image generation. This testbed features an extensive array of six LoRA categories, spanning two distinct visual styles: reality and anime. Our evaluation includes 480 diverse composition sets, each incorporating a varying number of LoRAs to comprehensively evaluate the efficacy of each proposed method. Given the lack of standardized automatic metrics for this novel task, we propose to employ GPT-4V (OpenAI, 2023a, b) as an evaluator, assessing both the quality of the images and the effectiveness of the compositions. Our empirical findings consistently demonstrate that both LoRA Switch and LoRA Composite substantially outperform the prevalent LoRA merging approach, particularly noticeable as the number of LoRAs in a composition increases. To further validate our results, we also conduct human evaluations, which reinforce our conclusions and affirm the efficacy of our automated evaluation framework. In addition, we provide a detailed analysis of the applicable scenarios for each method, as well as discuss the potential bias of using GPT-4V as an evaluator.
@@
(1) We introduce the first investigation of multi-LoRA composition from a decoding-centric perspective, proposing LoRA Switch and LoRA Composite. Our methods overcome existing constraints on the number of LoRAs that can be integrated, offering enhanced flexibility and improved quality in composable image generation.
(2) Our work establishes ComposLoRA, a comprehensive testbed tailored to this research area, featuring six varied categories of LoRAs and 480 composition sets. Addressing the absence of standardized metrics, we present an evaluator built upon GPT-4V, setting a new benchmark for assessing both image quality and compositional efficacy.
(3) Through extensive automatic and human evaluations, our findings reveal the superior performance of the proposed LoRA Switch and LoRA Composite compared to the prevalent LoRA merging approach. Additionally, we provide an in-depth analysis of different multi-composition methods and evaluation frameworks.
@@ 2.2 Multi-LoRA Composition through a Decoding-Centric Perspective
To address the above issues, we base our approach on the denoising process of diffusion models and investigate how to perform composition while maintaining the LoRA weights unchanged. This is specifically divided into two perspectives: in each denoising step, either activate only one LoRA or engage all LoRAs to guide the generation.
@@
To explore incorporating all LoRAs at each timestep without merging weight matrices, we propose LoRA Composite (LoRA-c), an approach grounded in the classifier-free guidance paradigm. This method involves calculating both unconditional and conditional score estimates for each LoRA individually at every denoising step. By aggregating these scores, the technique ensures balanced guidance throughout the image generation process, facilitating the cohesive integration of all elements represented by different LoRAs.
@@
Here, $w_{i}$ is a scalar weight allocated to each LoRA, intended to adjust the influence of the $i$-th LoRA. For our experimental setup, we set $w_{i}$ to 1, giving each LoRA equal importance. LoRA-c assures that every LoRA contributes effectively at each stage of the denoising process, addressing the potential issues of robustness and detail preservation that are commonly associated with merging LoRAs.
[…]
@@ 3.1 Experimental Setup
(1) Each LoRA should be robustly trained, ensuring it can accurately replicate the specific elements it represents when integrated independently;
(2) The elements represented by the LoRAs should cover a diverse range of categories and demonstrate adaptability across different image styles;
(3) When composed, LoRAs from different categories should be compatible, preventing any conflicts in the resulting image composition.
@@ 3. Object: two-handed burger, holding a huge burger
[[[[Evaluation criteria and format requirements]normal-]]]
@@
While existing metrics can calculate the alignment between text and images (Hessel et al., 2021; Ku et al., 2023), they fall short in assessing the intricacies of specific elements within an image and the quality of their composition. Recently, multimodal large language models like GPT-4V (OpenAI, 2023a, b) have significant progress and promise in various multimodal tasks, underscoring their potential in evaluating image generation task (Lin et al., 2023; Zhang et al., 2023b)s. In our study, we leverage GPT-4V’s capabilities to serve as an evaluator for composable image generation.
[…]
@@
For our experiments, we employ stable-diffusion-v1.5 (Rombach et al., 2022) as the backbone model. We utilize two specific checkpoints for our experiments: “Realistic_Vision_V5.1” [^2] https://huggingface.co/SG161222/Realistic_Vision_V5.1_noVAE. for realistic images and “Counterfeit-V2.5” [^3] https://huggingface.co/gsdf/Counterfeit-V2.5. for anime images, each fine-tuned to their respective styles. In the realistic style subset, we configure the model with 100 denoising steps, a guidance scale $s$ of 7, and set the image size to 1024x768, optimizing for superior image quality. For the anime style subset, the settings differ slightly with 200 denoising steps, a guidance scale $s$ of 10, and an image size of 512x512. The DPM-Solver++ (Lu et al., 2022a, b) is used as the scheduler in the generation process. The weight scale $w$ is consistently set at 0.8 for composing LoRAs within ComposLoRA. For the LoRA Switch approach, we apply a cycle with $\tau$ set to 5, meaning every 5 denoising steps activate the next LoRA in the sequence: character, clothing, style, background, then object. To ensure the reliability of our experimental results, we conduct image generation using three random seeds. All reported results in this paper represent the average evaluation scores across these three runs.
@@ 3.2 Results on ComposLoRA
We first present the comparative evaluation results obtained using GPT-4V. This evaluation involves scoring the performance of LoRA-s versus LoRA Merge, and LoRA-c versus LoRA Merge across two dimensions, as well as determining the winner based on these scores. Specific scores and win rates are illustrated in Figure LABEL:fig:result, leading to several key observations:
(1) Our proposed method consistently outperforms LoRA Merge across all configurations and in both dimensions, with the margin of superiority increasing as the number of LoRAs grows. For instance, as shown in Figure LABEL:fig:switch_scoring, the score advantage of LoRA Switch escalates from 0.04 with 2 LoRAs to 1.32 with 5 LoRAs. This trend aligns with the win rate observed in Figure LABEL:fig:switch_win_rate, where the win rate approaches 70% when composing 5 LoRAs.
(2) LoRA-S shows superior performance in composition quality, whereas LoRA-C excels in image quality. In scenarios involving 5 LoRAs and using LoRA Merge as a baseline, the win rate of LoRA-s in composition quality surpasses that of LoRA-c by 14% (69% vs. 55%). Conversely, for image quality, LoRA-c’s win rate is 10% higher than that of LoRA-s (56% vs. 46%).
(3) The task of compositional image generation remains highly challenging, especially as the number of elements to be composed increases. According to GPT-4V’s scoring, the average score for composing 2 LoRAs is above 8.5, but it sharply declines to around 6 for compositions involving 5 LoRAs. Hence, despite the considerable improvements our methods offer, there is still substantial room for further research in the field of compositional image generation.
Human Evaluation.
To complement our results, we conduct a human evaluation to assess the effectiveness of different methods and validate GPT-4V’s efficacy as an evaluator. Two graduate students rate 120 images on compositional and image quality using a 1-5 Likert scale: 1 signifies complete failure, 2-4 represent significant, moderate, and minor issues, respectively, while 5 denotes perfect execution. To ensure consistency, the annotators initially pilot-score 20 images to standardize their understanding of the criteria. The results, summarized in the upper section of Table 2, align with GPT-4V’s findings, confirming our methods outperform LoRA Merge — with LoRA Switch excelling in composition and LoRA Composite in image quality.
Furthermore, we analyze the Pearson correlations between human evaluations and scores derived from GPT-4V and CLIPScore (Hessel et al., 2021), with results presented in the lower section of Table 2. This comparison reveals that CLIPScore’s evaluations fall short in assessing specific compositional and quality aspects due to its inability to discern the nuanced features of each element. On the other hand, the GPT-4V-based evaluator we adopt shows substantially higher correlations with human judgments, affirming the validity of our evaluation framework.

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) is extensively utilized in text-to-image models for the accurate rendition of specific elements like distinct characters or unique styles in generated images. Nonetheless, existing methods face challenges in effectively composing multiple LoRAs, especially as the number of LoRAs to be integrated grows, thus hindering the creation of complex imagery. In this paper, we study multi-LoRA composition through a decoding-centric perspective. We present two training-free methods: LoRA Switch, which alternates between different LoRAs at each denoising step, and LoRA Composite, which simultaneously incorporates all LoRAs to guide more cohesive image synthesis. To evaluate the proposed approaches, we establish ComposLoRA, a new comprehensive testbed as part of this research. It features a diverse range of LoRA categories with 480 composition sets. Utilizing an evaluation framework based on GPT-4V, our findings demonstrate a clear improvement in performance with our methods over the prevalent baseline, particularly evident when increasing the number of LoRAs in a composition. The code, benchmarks, LoRA weights, and all evaluation details are available on our project website.
@@ 1 Introduction
In the dynamic realm of generative text-to-image models (Ho et al., 2020; Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; Ruiz et al., 2023; Sohn et al., 2023), the integration of Low-Rank Adaptation (LoRA) (Hu et al., 2022) stands out for its ability to fine-tune image synthesis with remarkable precision and minimal computational load. LoRA excels by specializing in one element — such as a specific character, a particular clothing, a unique style, or other distinct visual aspects — and being trained to produce diverse and accurate renditions of this element in generated images. For instance, users could customize their LoRA models to generate various images of themselves, achieving an array of personalized and realistic representations. The application of LoRA not only showcases its adaptability and precision in image generation but also opens new avenues in customized digital content creation, revolutionizing how users interact with and utilize generative text-to-image models for creating tailored visual content.
@@
However, an image typically embodies a mosaic of various elements, making compositionality key to controllable image generation (Tenenbaum, 2018; Huang et al., 2023b). In pursuit of this, the strategy of composing multiple LoRAs, each focused on a distinct element, emerges as a feasible approach for advanced customization. This technique enables the digitization of complex scenes, such as virtual try-ons, merging users with clothing in a realistic fashion, or urban landscapes where users interact with meticulously designed city elements. Prior investigations into multi-LoRA compositions have explored the context of pre-trained language models (Zhang et al., 2023a; Huang et al., 2023a) or stable diffusion models (Ryu, 2023; Shah et al., 2023). These studies aim to merge multiple LoRA models to synthesize a new LoRA model by training coefficient matrices (Huang et al., 2023a; Shah et al., 2023; Wu et al., 2024) or through the direct addition or subtraction of LoRA weights (Ryu, 2023; Zhang et al., 2023a). Nevertheless, these approaches centered on weight manipulation could destabilize the merging process as the number of LoRAs grows (Huang et al., 2023a) and also overlook the interaction between LoRA models and base models. This oversight becomes particularly critical in diffusion models, which depend on sequential denoising steps for image generation. Ignoring the interplay between LoRAs and these steps can result in misalignments in the generative process, as shown in Figure 1, where a merged LoRA model fails to preserve the full complexity of all desired elements, leading to distorted or unrealistic images.
@@
Figure 1: Multi-LoRA composition techniques effectively blend different elements such as characters, clothing, and objects into a cohesive image. Unlike the conventional LoRA Merge approach (Ryu, 2023), which can lead to detail loss and image distortion as more LoRAs are added, our methods retain the accuracy of each element and the overall image quality.
@@
In this paper, we delve into multi-LoRA composition from a decoding-centric perspective, keeping all LoRA weights intact. We present two training-free approaches that utilize either one or all LoRAs at each decoding step to facilitate compositional image synthesis. Our first approach, LoRA Switch, operates by selectively activating a single LoRA during each denoising step, with a rotation among multiple LoRAs throughout the generation process. For instance, in a virtual try-on scenario, LoRA Switch alternates between a character LoRA and a clothing LoRA at successive denoising steps, thereby ensuring that each element is rendered with precision and clarity. In parallel, we propose LoRA Composite, a technique that draws inspiration from classifier-free guidance (Ho & Salimans, 2022). It involves calculating unconditional and conditional score estimates derived from each respective LoRA at every denoising step. These scores are then averaged to provide balanced guidance for image generation, ensuring a comprehensive incorporation of all elements. Furthermore, by bypassing the manipulation on the weight matrix but directly influencing the diffusion process, both methods allow for the integration of any number of LoRAs and overcome the limitations of recent studies that typically merge only two LoRAs (Shah et al., 2023).
@@
Experimentally, we introduce ComposLoRA, the first testbed specifically designed for LoRA-based composable image generation. This testbed features an extensive array of six LoRA categories, spanning two distinct visual styles: reality and anime. Our evaluation includes 480 diverse composition sets, each incorporating a varying number of LoRAs to comprehensively evaluate the efficacy of each proposed method. Given the lack of standardized automatic metrics for this novel task, we propose to employ GPT-4V (OpenAI, 2023a; b) as an evaluator, assessing both the quality of the images and the effectiveness of the compositions. Our empirical findings consistently demonstrate that both LoRA Switch and LoRA Composite substantially outperform the prevalent LoRA merging approach, particularly noticeable as the number of LoRAs in a composition increases. To further validate our results, we also conduct human evaluations, which reinforce our conclusions and affirm the efficacy of our automated evaluation framework. In addition, we provide a detailed analysis of the applicable scenarios for each method, as well as discuss the potential bias of using GPT-4V as an evaluator.
@@
  • We introduce the first investigation of multi-LoRA composition from a decoding-centric perspective, proposing LoRA Switch and LoRA Composite. Our methods overcome existing constraints on the number of LoRAs that can be integrated, offering enhanced flexibility and improved quality in composable image generation.
  • Our work establishes ComposLoRA, a comprehensive testbed tailored to this research area, featuring six varied categories of LoRAs and 480 composition sets. Addressing the absence of standardized metrics, we present an evaluator built upon GPT-4V, setting a new benchmark for assessing both image quality and compositional efficacy.
  • Through extensive automatic and human evaluations, our findings reveal the superior performance of the proposed methods compared to the prevalent LoRA merging approach. Additionally, we provide an in-depth analysis of different multi-composition methods and evaluation frameworks.
2 Related Work
2.1 Composable Text-to-Image Generation
Composable image generation, a key aspect of digital content customization, involves creating images that adhere to a set of pre-defined specifications (Liu et al., 2023). Existing research in this domain primarily focuses on the following approaches: enhancing compositionality with scene graphs or layouts (Johnson et al., 2018; Yang et al., 2022; Gafni et al., 2022), modifying the generative process of diffusion models to align with the underlying specifications (Feng et al., 2023; Huang et al., 2023c; b), multi-concept customization (Kumari et al., 2023; Han et al., 2023; Gu et al., 2023; Kwon et al., 2024; Kong et al., 2024), or composing a series of independent models that enforce desired constraints (Du et al., 2020; Liu et al., 2021; Nie et al., 2021; Liu et al., 2022; Li et al., 2023; Du et al., 2023).
However, these methods typically operate at the concept level, where generative models excel in creating images based on broader categories or general concepts. For example, a model might be prompted to generate an image of “a woman wearing a dress”, and can adeptly accommodate variations in the textual description, such as changing the color of the dress. Yet, they struggle to accurately render specific, user-defined elements, like lesser-known characters or unique dress styles. Another line of work that can compose user-defined objects into images (Huang et al., 2023c; Ruiz et al., 2023). However, these methods require extensive fine-tuning and do not perform well on multiple objects. Therefore, we introduce learning-free instance-level composition approaches utilizing LoRA, enabling the precise assembly of user-specified elements in image generation.
2.2 LoRA-based Manipulations
@@ 3.2 Multi-LoRA Composition through a Decoding-Centric Perspective
To address the above issues, we base our approach on the denoising process and investigate how to perform composition while maintaining the LoRA weights unchanged. This is specifically divided into two perspectives: in each denoising step, either activate only one LoRA or engage all LoRAs to guide the generation.
@@
To explore incorporating all LoRAs at each timestep without merging weight matrices, we propose LoRA Composite (LoRA-c), an approach grounded in the Classifier-Free Guidance paradigm. Previous research has primarily focused on modifying CFG to enable diffusion models to emphasize textual concepts (Liu et al., 2022; Du et al., 2023; Sohn et al., 2023). In contrast, our method extends this by enabling CFG to condition on LoRAs, facilitating the generation of images that reflect specific elements or instances rather than abstract concepts. LoRA-c involves calculating both unconditional and conditional score estimates for each LoRA individually at every denoising step. By aggregating these scores, the technique ensures balanced guidance throughout the image generation process, facilitating the cohesive integration of all elements represented by different LoRAs.
@@
Here, $w_{i}$ is a scalar weight allocated to each LoRA, intended to adjust the influence of the $i$-th LoRA. In this paper, we set $w_{i}$ to 1, giving each LoRA equal importance. LoRA-c assures that every LoRA contributes effectively at each stage of the denoising process, addressing the potential issues of robustness and detail preservation that are commonly associated with merging LoRAs.
Overall, we are the first to adopt a decoding-centric perspective in multi-LoRA composition, steering clear of the instability inherent in weight manipulation on LoRAs. Our study introduces two training-free methods for activating either one or all LoRAs at each denoising step, with their comparative analysis presented in Sections §$\S$4.2 and §$\S$4.3.1.
@@ 4.1 Experimental Setup
  • Each LoRA should be robustly trained, ensuring it can accurately replicate the specific elements it represents when integrated independently;
  • The elements represented by the LoRAs should cover a diverse range of categories and demonstrate adaptability across different image styles;
  • When composed, LoRAs from different categories should be compatible, preventing any conflicts in the resulting image composition.
@@ 3. Object: two-handed burger, holding a huge burger
[[[[Evaluation criteria and format requirements]]]]
@@
While existing metrics can calculate the alignment between text and images (Hessel et al., 2021; Ku et al., 2023), they fall short in assessing the intricacies of specific elements within an image and the quality of their composition. Recently, multimodal large language models like GPT-4V (OpenAI, 2023a; b) have significant progress and promise in various multimodal tasks, underscoring their potential in evaluating image generation tasks (Lin et al., 2023; Zhang et al., 2023b). In our study, we leverage GPT-4V’s capabilities to serve as an evaluator for composable image generation.
Specifically, we employ a comparative evaluation method, utilizing GPT-4V to rate generated images across two dimensions: composition quality and image quality. We utilize a 0 to 10 scoring scale, with higher scores indicating superior quality. GPT-4V is provided with a prompt that includes the essential features of the elements to be composed, the criteria for scoring in the two dimensions, and the format for the expected output. The complete evaluation prompts and results are available in Tables 7 and 8 in Appendix. This experimental setup allows us to compare the efficacy of each of the two proposed methods against the LoRA Merge approach. Additionally, we examine how GPT-4V-based scoring aligns with human judgment in Section §$\S$4.2 and explore the potential biases of using it as an evaluator in Section §$\S$4.3.3.
@@
For our experiments, we employ stable-diffusion-v1.5 (Rombach et al., 2022) as the backbone model. We utilize two specific checkpoints for our experiments: “Realistic_Vision_V5.1” [^2] https://huggingface.co/SG161222/Realistic_Vision_V5.1_noVAE. for realistic images and “Counterfeit-V2.5” [^3] https://huggingface.co/gsdf/Counterfeit-V2.5. for anime images, each fine-tuned to their respective styles. In the realistic style subset, we configure the model with 100 denoising steps, a guidance scale $s$ of 7, and set the image size to 1024x768, optimizing for superior image quality. For the anime style subset, the settings differ slightly with 200 denoising steps, a guidance scale $s$ of 10, and an image size of 512x512. The DPM-Solver++ (Lu et al., 2022a; b) is used as the scheduler in the generation process. The weight scale $w$ is consistently set at 0.8 for composing LoRAs within ComposLoRA. For the LoRA Switch approach, we apply a cycle with $\tau$ set to 5, meaning every 5 denoising steps activate the next LoRA in the sequence: character, clothing, style, background, then object. Since the proposed methods do not require additional training, all experiments are conducted on a single A6000 GPU. To ensure the reliability of our experimental results, we conduct image generation using three random seeds. All reported results in this paper represent the average evaluation scores across these three runs.
@@ 4.2 Results on ComposLoRA
We first present the comparative evaluation results using GPT-4V. This evaluation involves scoring the performance of LoRA-s versus LoRA Merge, and LoRA-c versus LoRA Merge across two dimensions, as well as determining the winner based on these scores. Specific scores and win rates are illustrated in Figure 3, leading to several key observations:
(a) LoRA Switch vs. LoRA Merge (Scoring)
(b) LoRA Composite vs. LoRA Merge (Scoring)
(c) LoRA Switch vs. LoRA Merge (Win Rate)
(d) LoRA Composite vs. LoRA Merge (Win Rate)
Figure 3: Results of comparative evaluation on ComposLoRA using GPT-4V.
  • Our proposed method consistently outperforms LoRA Merge across all configurations and in both dimensions, with the margin of superiority increasing as the number of LoRAs grows. For instance, as shown in Figure 3(a), the score advantage of LoRA Switch escalates from 0.04 with 2 LoRAs to 1.32 with 5 LoRAs. This trend aligns with the win rate observed in Figure 3(c), where the win rate approaches 70% when composing 5 LoRAs.
  • LoRA-S shows superior performance in composition quality, whereas LoRA-C excels in image quality. In scenarios involving 5 LoRAs and using LoRA Merge as a baseline, the win rate of LoRA-s in composition quality surpasses that of LoRA-c by 14% (69% vs. 55%). Conversely, for image quality, LoRA-c’s win rate is 10% higher than that of LoRA-s (56% vs. 46%).
  • The task of compositional image generation remains highly challenging, especially as the number of elements to be composed increases. According to GPT-4V’s scoring, the average score for composing 2 LoRAs is above 8.5, but it sharply declines to around 6 for compositions involving 5 LoRAs. Hence, despite the considerable improvements our methods offer, there is still substantial room for further research in the field of compositional image generation.
@@
Human Evaluation.
To complement our results, we conduct a human evaluation to assess the effectiveness of different methods and validate the efficacy of the evaluators.
Two graduate students rate 120 images on compositional and image quality using a 1-5 Likert scale: 1 signifies complete failure, 2-4 represents significant, moderate, and minor issues, respectively, while 5 denotes perfect execution. To ensure consistency, the annotators initially pilot-score 20 images to standardize their understanding of the criteria. The results, summarized in the upper section of Table 2, align with GPT-4V’s findings, confirming our methods outperform LoRA Merge — with LoRA Switch excelling in composition and LoRA Composite in image quality.
Furthermore, we analyze the Pearson correlations between human evaluations and scores derived from GPT-4V and CLIPScore (Hessel et al., 2021), with results presented in the lower section of Table 2. This comparison reveals that CLIPScore’s evaluations fall short in assessing specific compositional and quality aspects due to its inability to discern the nuanced features of each element. In contrast, the evaluator we adopt shows substantially higher correlations with human judgments, affirming the validity of our evaluation framework.
4.3 Analysis
To enhance our understanding of the proposed methods, we further investigate the following questions: